OpenStudy (jkbo):

What is the next term in the sequence below? 1, 3, 5, 8, 11, 15, 19, 24, ... (1 point) 30 27 29 25

OpenStudy (diyadiya):

1 + 2 =3 3 + 2 =5 5+3=8 8+3=11 11+4=15 15+4=19 19+5=24 so 24+5=29

OpenStudy (jkbo):

thanks